# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ 5 Results
The Gaussian initial conditions used for most simulations are in approximate vertical equilibrium but highly unstable to the Parker instability and MRI. The system thus rapidly reorganizes over the course of $2$-$3$ orbits, transporting significant magnetic flux to larger $z$ where it can escape out of the top and bottom boundaries, thereby organizing into a lower-$\beta$ corona if the vertical domain is sufficiently large. During this reorganization phase, around half of the initial azimuthal magnetic flux is lost from the domain if starting from Gaussian initial conditions (this fraction seems to be independent of $\beta$), while for the quasi-Lorentzian initial conditions of Eq. 5 the fraction lost is lower because the initial state is rather similar to the final one. As per points 1-2 in §2, we find that two very different quasi-steady states result after it settles, the transition between which seems to be controlled by the midplane $\beta$ following the transient reorganization phase. If too little flux remains after this phase, it all escapes. Alternatively, if the system remains sufficiently magnetized, the flux’s escape is balanced by its regeneration and is maintained in steady state. For constant-$\beta$ Gaussian initial conditions, we find that the tipping point occurs for an initial $\beta\simeq 0.1$: higher-$\beta$ initial conditions transition into the high-$\beta$ state, lower-$\beta$ ones transition into the low-$\beta$ state, and those with $\beta=0.1$ are sensitive the details, as described below.
@@
We illustrate this transition in Fig. 1, which shows so-called “butterfly diagrams” of the time evolution of $\overline{\bm{B}}(z)$ (the $x,y$ mean of $\bm{B}$). The top panels show cases that transition into the high-$\beta$ state, which is characterized by dynamo cycles that quasi-periodically reverse the azimuthal field, in good agreement with previous work (Simon et al., 2012; Salvesen et al., 2016a). The three cases shown, $\beta$1, $\beta$0.3, and $\beta$0.1 have the same initial $H_{\beta}$ but successively decreasing initial $\beta$, and thus $H_{\rm th}$, in the Gaussian initial conditions (Eq. 3). Each case loses nearly all of its initial midplane magnetic flux to yield a thermally dominated midplane. The dynamo cycles are well formed and coherent, with a similar period of $\simeq 10$ orbits in each case, showing that the cycle period is independent of the sound speed (as expected since similar physics seems to occur in homogenous and/or incompressible models; Lesur & Ogilvie, 2008b; Shi et al., 2016, e.g.,).
@@
The lower panels show results initialized with more flux, which instead sustain the low-$\beta$ state. The first, $\beta$R0.1 is identical to $\beta$0.1 except that it is also initialized with a mean radial field with $\beta_{x}=100$ (the sign is such that $B_{x}B_{y}<0$ as required to strengthen the initial $B_{y}$). The dynamics are very different to $\beta$0.1 — there are no cycles and the system maintains a much stronger $B_{y}$ in the midplane. Lower-initial-$\beta$ simulations with smaller $H_{\rm th}$ ($\beta$0.05 and $\beta$0.01) reach a very similar state, albeit with a narrower midplane peak because this width is ultimately set by $H_{\rm th}$ (see below). Changes to $\overline{B_{y}}$ in time are small amplitude, slow, and seemingly random, including periods in which the midplane field strengthens, thus showing that these simulations are not simply taking longer to lose their flux. Initial conditions with $\beta=0.1$ seem to lie on boundary between reaching the low- and high-$\beta$ states. This is evidenced by the fact that $\beta$0.1 loses its flux significantly more slowly than $\beta$1 and $\beta$0.3, and a lower-resolution version of the same setup ($\beta 0.1$-mr) reaches the low-$\beta$ state. This latter outcome appears to be a consequence of the expulsion of the flux being somewhat less efficient at lower resolution — when initialized with an additional very weak ($\beta_{x}=1000$) radial field that weakens the initial $B_{y}$ (i.e., $B_{x}B_{y}>0$), the same setup instead transitions into the high-$\beta$ state (simulation $\beta$nR0.1-mr). The simulation $\beta$0.1-mr has been run over a significantly longer time, around $100$ orbits, in order to test for sudden or long-term changes to the low-$\beta$ state. Its lack of significant change over this time suggests that the system can maintain the low-$\beta$ state indefinitely.
@@
$\beta$R0.1-noshear 0.1 100 $\infty$ 0.43 $336^{3}$ (10,48) Power-law no shear High-$\beta$0.014
$\beta$R0.1-norot 0.1 100 $\infty$ 0.43 $336^{3}$ (10,48) Power-law no rotation High-$\beta$ 0.014
$\beta$0.01-noshear 0.01 $\infty$ $\infty$ 0.14 $336^{3}$ (20,34) Outflow no shear High-$\beta$ 0.17
$\beta$0.01-norot 0.01 $\infty$ $\infty$ 0.14 $336^{3}$ (20,42) Outflow no rotation High-$\beta$ 0.14
@@
$\beta$R0.1-NVF2-norot 0.1 $100$ $100$ 0.43 $336^{3}$ (5,27) Power-law no rotation High-$\beta$ 0.17
Table 1: A list of the main simulations analyzed in this work with relevant parameters, initial conditions, and other properties. The simulation Name is chosen based on the initial $\beta$ in a Gaussian equilibrium (3) with $H_{\beta}=\sqrt{2}$, thus effectively providing a correspondence with the thermal scale height in box units (column 4). The seventh column lists the final time and the “steady state” time $t_{\rm av}$, with averages performed for $t\in(t_{\rm av},t_{\rm fin})$. Boundary conditions (“BCs”) are described in §3.2. “Outcome” lists whether the simulation reaches the low- or high-$\beta$ state, which is determined by its time-averaged profile and $\alpha$ (final column). Other simulations used for numerical tests are listed in table 2.
@@
The aforementioned behavior is illustrated more quantitatively in Figs. 3 and 4, which show, respectively, time-averaged density and magnetic-pressure profiles, and the time evolution of key quantities. The former shows how high-$\beta$ profiles are thermally dominated out to $\simeq 2H_{\rm th}$, with a lower-$\beta$ corona above this (note that $P_{B}$ here is defined using the mean fields as $P_{B}=\overline{B_{y}}^{2}/8\pi$ averaged over time; when fluctuations are also included in $P_{B}$, $\beta\approx 40$ in the midplane) [^3] The mean density profile all three cases shown here approaches artificial density floor of $10^{-4}$ at the $z$ boundaries. This may artificially flatten the density profile and, as a consequence, $P_{B}$ at large $z$, thus also reducing high-$z$ turbulent fluctuations due to the weaker Parker instability.. The low-$\beta$ state is magnetically dominated everywhere, though increasingly so at large $z$, where the profiles are approximate power laws with an index $a\approx 2.75$. The lowest-$\beta$ initial conditions ($\beta$0.01 and $\beta$0.001) also retain $\beta\ll 1$ midplanes in steady state, although we argue below that this is likely a consequence of not adequately resolving $H_{\rm th}$, which sets the “curvature” of $\rho$ for $z\lesssim H_{\rm th}$ (see dotted lines for $\beta 0.1$R). The small dip in $P_{B}$ in the midplane is a persistent feature when properly resolved; this is a consequence of the pressure support being thermal rather than magnetic at the midplane. This also implies that the Parker instability is quenched near the midplane since $zdP_{B}/dz>0$.
@@
Figure 4: Time evolution of various important quantities from the simulations shown in Fig. 1, as well as several others that illustrate important features. The top panel shows the total accretion stress $\alpha=\langle{\rho u_{x}u_{y}-B_{x}B_{y}}\rangle/\langle{\rho c_{s}^{2}}\rangle$, and the second and third panels show the midplane $\beta$ and magnetic flux $\langle{B_{y}}\rangle_{\rm th}/\langle{B_{y}(t=0)}\rangle_{\rm th}$, respectively, where the midplane is defined here as the region $|z|\leq H_{\rm th}$ (the thin black line on $\langle{B_{y}}_{\rm th}\rangle/\langle{B_{y}(t=0)}_{\rm th}\rangle$ shows $2\pi\Omega^{-1}d\ln{\langle{B_{y}}\rangle_{\rm th}}/dt\approx 1/6$ for later discussion). We see clearly bimodal behavior in all quantities for various initial conditions and resolutions, although which state it reaches can have dependence on numerical parameters for cases near the transition (cf. $\beta$0.1 and $\beta$0.1-mr). The initial $\beta_{\rm mid}$ is higher for the cases starting from Lorentzian-like initial conditions because this equilibrium already has a strongly magnetized corona at larger $z$. The bottom panel shows the same quantities for $\beta$0.1-m over its full time duration of ${\simeq}100$ orbits. We do not observe any long-term or sudden changes to the low-$\beta$ state.
@@
Figure 4 shows the time evolution of the total accretion stress $\alpha\equiv(\langle{\rho u_{x}u_{y}}\rangle-\langle{B_{x}B_{y}}\rangle)/\langle{P_{\rm th}}\rangle$, the midplane $\beta$ including fluctuations ($\beta_{\rm mid}\equiv 8\pi c_{s}^{2}\langle{\rho}\rangle_{\rm th}/\langle{B^{2}}\rangle_{\rm th}$), and the midplane $\overline{B_{y}}$ for the simulations shown in Fig. 1 as well as several other illustrative examples. Most importantly, we see a large “gap” between the low-$\beta$ and high-$\beta$ states for each quantity. As well as clearly illustrating the bimodality in $\beta$ between the two states, these show more concretely that (i) the low-$\beta$ state supports much higher accretion rates and (ii) the low-$\beta$ state retains much more of the midplane magnetic flux from the initial conditions, whereas nearly all of the flux ($\gtrsim 95\%$) is lost in the evolution towards the high-$\beta$ state. It is also clear how the $\beta=0.1$, $H_{\beta}=\sqrt{2}$ initial conditions ($\beta$0.1, $\beta$R0.1, $\beta$0.1-mr, and $\beta$nR0.1-mr) are sensitive to slight changes in resolution and initial conditions: $\beta$0.1 transitions to the high-$\beta$ state rather slowly, while $\beta$R0.1 stays at low $\beta$ due to the extra initial contribution from $B_{x}$; $\beta$0.1-mr (at lower resolution) evolves back to lower $\beta$ and higher $\alpha$ after initially evolving similarly to $\beta$0.1, but adding a very weak radial field with $B_{x}B_{y}>0$ ($\beta$nR0.1-mr) causes it to transition to the high-$\beta$ state (note that this $B_{r}$ with $\beta_{x}=1000$, is weaker than that which occurs self-consistently in the low-$\beta$ saturated state). We also show two cases that start instead from the quasi-Lorentzian initial conditions of Eq. 5. Because these profiles are already magnetically dominated at large $z$, they evolve less violently over the initial phases. Nonetheless, the bimodal behavior persists, with the lower-$\beta$ initial condition maintaining the strongly magnetized state and $\overline{B_{y}}$ increasing, while the higher-$\beta$ initial condition loses nearly all of its midplane flux. [^4] At $336^{3}$ resolution and $H_{\rm th}=0.43$, the high-$\beta$ state only sometimes produces well-formed cyclic behavior. Given the low effective resolution of the thermal midplane, and previous results documenting the resolution and box-size dependence of MRI turbulence (e.g., Davis et al., 2010; Ryan et al., 2017), this is not unexpected.
@@
Figure 5: Spatial profile of the various components of the accretion stress in the low-$\beta$ state simulation $\beta$R0.1. We split $\alpha(z)$ into $\alpha_{K}\equiv\overline{u_{x}u_{y}}$, $\alpha_{\delta B}=-\overline{\delta B_{x}\delta B_{y}}$, and $\alpha_{\overline{B}}=-{\overline{B_{x}}\,\overline{B_{y}}}$, each normalized by $\langle{\rho}\rangle_{\rm th}c_{s}^{2}$. Magnetic contributions are dominant, with the total stress following a similar profile to $P_{B}$ except right in the midplane.
@@ 5.2 Turbulence and transport at low $\beta$
The turbulence profiles in the low-$\beta$ state are shown in Figs. 5 and 6. Figure 5 shows $\overline{\alpha}(z)$ in $\beta$R0.1, splitting the total into its contributions from velocity $\alpha_{K}=\overline{\rho u_{x}u_{y}}/\langle{P_{\rm th}}\rangle_{\rm th}$, magnetic fluctuations $\alpha_{\delta B}=-\overline{\delta B_{x}\delta B_{y}}/\langle{P_{\rm th}}\rangle_{\rm th}$, and the mean magnetic field $\alpha_{\overline{B}}=-\overline{B_{x}}\overline{B_{y}}/\langle{P_{\rm th}}\rangle_{\rm th}$. The total $\overline{\alpha}(z)$ profile is similar to that of the equilibrium-field $P_{B}$ (dotted line), aside from in the midplane, where there is a peak in the turbulent contributions and a drop off in the mean contribution. We interpret this change within $|z|\lesssim H_{\rm th}$ as being a consequence of the MRI taking over compared to the Parker instability, which dominates at larger $z$. Both sources of turbulence allow accretion ($\alpha>0$), but the fluctuations in the midplane region have a larger relative correlation between the $x$ and $y$ components (i.e., more efficient accretion for a given turbulence amplitude) as well as a stronger relative contribution from magnetic, compared to velocity fluctuations.
@@
The thin black line in both panels illustrates a simple model of the turbulence, based on driving by a combination of the MRI and Parker instability (see also Johansen & Levin 2008). For both instabilities, we apply the standard reasoning that the turbulence saturates when the nonlinear turnover time $\tau_{\rm nl}\sim\ell/\delta u$ at the outer scale $\ell$ balances the instability’s growth rate. In regions far above the midplane, where the profiles of $B_{y}$, $\rho$, and $v_{\rm A}$ are all power laws, it is most natural to take $\ell\propto z$, such that the turbulence remains self-similar with $z$. Another possible choice, the lengthscale of the fastest growing mode, would scale in the same way for both the Parker and MRI: for Parker because it grows fastest at the largest available vertical scales (the global scales, $\propto z$), and for MRI because it grows fastest where $n\sim\Omega^{-1}v_{\rm A}/\ell\sim 1$ and $v_{\rm A}\propto z$ in such regions (see §3.1 and inset of Fig. 6). [^5] Similar arguments applied to the high-$\beta$ state would suggest that we should take $\ell\sim v_{\rm A}/\Omega$ of the fastest growing mode as the outer scale, where the $v_{\rm A}$ is that of the self-induced mean azimuthal field that can drive the MRI. This scale is very small, around $\ell\sim 4\times 10^{-2}H_{\rm th}$ for the $\beta$1 simulation (effectively the grid scale), in reasonable agreement with the morphology seen in Fig. 2. If $\ell$ is indeed grid limited, then since $\delta u\propto\ell$ (assuming nearly constant growth rate), we would expect the turbulence amplitude to decrease with increasing resolution (complicated, perhaps, by the fact that $\overline{v_{\rm A}}$ may itself depend on $\delta u$). This is consistent with the results of Ryan et al. (2017) who found that the turbulence decreased its amplitude and scale with increasing resolution in the high-$\beta$ state, even at extremely high resolutions. In contrast, near the midplane, since the curvature of $\rho$ is set by $H_{\rm th}$, we require $\ell\lesssim H_{\rm th}$. The Parker instability will not operate here because $B_{y}$ is nearly flat, while the scale at which the MRI growth rate peaks ($n\approx 1$ in Eq. 6 with $v_{\rm A}\sim c_{s}$) is $\ell\sim v_{\rm A}/\Omega\sim H_{\rm th}$, suggesting $\ell\sim H_{\rm th}$ is a reasonable choice (this is also plausibly consistent with the morphology seen in Fig. 2 for $\beta$R0.1).
@@
Figure 7: Testing different vertical boundary conditions, numerical options, and initial conditions (see table 1). As in Fig. 3, solid lines show $P_{B}$ and dashed lines show $P_{\rm th}$.

### v3

@@ 5 Results
The Gaussian initial conditions used for most simulations are in approximate vertical equilibrium but highly unstable to the Parker instability and MRI. The system thus rapidly reorganizes over the course of $2$-$3$ orbits, transporting significant magnetic flux to larger $z$ where it can escape out of the top and bottom boundaries, thereby organizing into a lower-$\beta$ corona if the vertical domain is sufficiently large. During this reorganization phase, around half of the initial azimuthal magnetic flux is lost from the domain if starting from Gaussian initial conditions (this fraction seems to be independent of $\beta$), while for the quasi-Lorentzian initial conditions of Eq. 5 the fraction lost is lower because the initial state is rather similar to the final one. As per points 1-2 in §2, we find that two very different quasi-steady states result after it settles, the transition between which seems to be controlled by the midplane $\beta$ following the transient reorganization phase. If too little flux remains after this phase, it all escapes. Alternatively, if the system remains sufficiently magnetized, the flux’s escape is balanced by its regeneration and is maintained in steady state. For constant-$\beta$ Gaussian initial conditions, we find that the tipping point occurs for an initial $\beta\simeq 0.1$: higher-$\beta$ initial conditions transition into the high-$\beta$ state, lower-$\beta$ ones transition into the low-$\beta$ state, and those with $\beta=0.1$ are sensitive to the details, as described below. Note, however, that this $\beta\simeq 0.1$ threshold is specific to the choice of the constant-$\beta$ Gaussian: for the Lorentzian-like form (5), the system transitions into the low-$\beta$ state from larger initial midplane $\beta$ (e.g., $\approx 0.35$ for $\beta$0.1-LorH) because the vertical equilibrium starts closer to its steady state structure.
@@
We illustrate this transition in Fig. 2, which shows so-called “butterfly diagrams” of the time evolution of $\overline{\bm{B}}(z)$ (the $x,y$ mean of $\bm{B}$). The top panels show cases that transition into the high-$\beta$ state, which is characterized by dynamo cycles that quasi-periodically reverse the azimuthal field, in good agreement with previous work (Simon et al., 2012; Salvesen et al., 2016a). The three cases shown, $\beta$1, $\beta$0.3, and $\beta$0.1 have the same initial $H_{\beta}$ but successively decreasing initial $\beta$, and thus $H_{\rm th}$, in the Gaussian initial conditions (Eq. 3). Each case loses nearly all of its initial midplane magnetic flux to yield a thermally dominated midplane. The dynamo cycles are well formed and coherent, each with a similar period of $\simeq 10$ orbits, showing that the cycle period is independent of the sound speed (as expected since similar physics seems to occur in homogenous and/or incompressible models; Lesur & Ogilvie, 2008b; Shi et al., 2016, e.g.,).
@@
The lower panels show results initialized with more flux, which instead sustain the low-$\beta$ state. The first, $\beta$R0.1 is identical to $\beta$0.1 except that it is also initialized with a mean radial field with $\beta_{x}=100$ (the sign is such that $B_{x}B_{y}<0$ as required to strengthen the initial $B_{y}$). The dynamics are very different to $\beta$0.1 — there are no cycles and the system maintains a much stronger $B_{y}$ in the midplane. Lower-initial-$\beta$ simulations with smaller $H_{\rm th}$ ($\beta$0.05 and $\beta$0.01) reach a very similar state, albeit with a narrower midplane peak because this width is ultimately set by $H_{\rm th}$ (see below). Changes to $\overline{B_{y}}$ in time are small amplitude, slow, and seemingly random. There also occurs times over which the midplane field strengthens, showing that these simulations are not simply taking longer to lose their flux. Initial conditions with $\beta=0.1$ seem to lie on the boundary between reaching the low- and high-$\beta$ states. This is evidenced by the fact that $\beta$0.1 loses its flux significantly more slowly than $\beta$1 and $\beta$0.3, and a coarser-resolution version of the same setup ($\beta 0.1$-mr) reaches the low-$\beta$ state. This latter outcome appears to be a consequence of the expulsion of the flux being somewhat less efficient at coarser resolution: when initialized with an additional very weak ($\beta_{x}=1000$) radial field that weakens the initial $B_{y}$ (i.e., $B_{x}B_{y}>0$), the same setup instead transitions into the high-$\beta$ state (simulation $\beta$nR0.1-mr). The simulation $\beta$0.1-mr has been run over a significantly longer time, around $100$ orbits, in order to test for sudden or long-term changes to the low-$\beta$ state. Its lack of significant change over this time suggests that the system can maintain the low-$\beta$ state indefinitely.
@@
Table 1: A list of the main simulations analyzed in this work with relevant parameters, initial conditions, and other properties. The simulation Name is chosen based on the initial $\beta$ in a Gaussian equilibrium (3) with $H_{\beta}=\sqrt{2}$, thus effectively providing a correspondence with the thermal scale height in box units (column 5). The seventh column lists the final time and the “steady state” time $t_{\rm av}$, with averages performed for $t\in(t_{\rm av},t_{\rm fin})$. Boundary conditions (“BCs”) are described in §3.2. “Outcome” lists whether the simulation reaches the low- or high-$\beta$ state, which is determined by its time-averaged profile and $\alpha$ (final column). Other simulations used for numerical tests are listed in table 2.
@@
The aforementioned behavior is illustrated more quantitatively in Figs. 4 and 5, which show, respectively, time-averaged density and magnetic-pressure profiles, and the time evolution of key quantities. The former shows how high-$\beta$ profiles are thermally dominated out to $\simeq 2H_{\rm th}$, with a lower-$\beta$ corona above this (note that $P_{B}$ here is defined using the mean fields as $P_{B}=\overline{B_{y}}^{2}/8\pi$ averaged over time; when fluctuations are also included in $P_{B}$, $\beta\approx 40$ in the midplane). [^5] The mean density profile in all three cases shown here approaches the artificial density floor of $10^{-4}$ at the $z$ boundaries. This may artificially flatten the density profile and, as a consequence, $P_{B}$ at large $z$, thus also reducing high-$z$ turbulent fluctuations due to the weaker Parker instability. The low-$\beta$ state is magnetically dominated everywhere, though increasingly so at large $z$, where the profiles are approximate power laws with an index $a\approx 2.75$ in Eq. 4. The lowest-$\beta$ initial conditions ($\beta$0.01 and $\beta$0.001) also retain $\beta\ll 1$ midplanes in steady state, although we argue below that this is likely a consequence of not adequately resolving $H_{\rm th}$, which sets the shape of $\rho$ for $z\lesssim H_{\rm th}$ (see dotted lines for $\beta$R0.1). The small dip in $P_{B}$ in the midplane is a persistent feature when properly resolved; this is a consequence of the pressure support being thermal rather than magnetic at the midplane and also implies that the Parker instability is quenched there since $zdP_{B}/dz>0$. By solving Eq. 2 for $\overline{B_{y}}$ from the measured $\overline{\rho}$, then comparing this to the measured $\overline{B_{y}}$ (or vice versa from $\overline{B_{y}}$ to $\overline{\rho}$), we find that the mean low-$\beta$ profiles in Fig. 4 are almost exactly equilibria (not shown). This is expected because the turbulence remains sub-Alvénic, justifying directly the magneto-static approach of §3.1.
@@
Figure 5: Time evolution of various important quantities from the simulations shown in Fig. 2, as well as several others that illustrate important features. The top panel shows the total accretion stress $\alpha$, and the second and third panels show the midplane $\beta$ and magnetic flux $\langle{B_{y}}\rangle_{\rm th}/\langle{B_{y}(t=0)}\rangle_{\rm th}$, respectively, where the midplane is defined here as the region $|z|\leq H_{\rm th}$ (the thin black line on $\langle{B_{y}}_{\rm th}\rangle/\langle{B_{y}(t=0)}_{\rm th}\rangle$ shows $2\pi\Omega^{-1}d\ln{\langle{B_{y}}\rangle_{\rm th}}/dt\approx 1/6$ for later discussion). We see clearly bimodal behavior in all quantities for various initial conditions and resolutions, although which state it reaches can have dependence on numerical parameters for cases near the transition (cf. $\beta$0.1 and $\beta$0.1-mr). The initial $\beta_{\rm mid}$ is higher for the cases starting from Lorentzian-like initial conditions because this equilibrium already has a strongly magnetized corona at larger $z$. The bottom panel shows the same quantities for $\beta$0.1-mr over its full time duration of ${\simeq}100$ orbits. We do not observe any long-term or sudden changes to the low-$\beta$ state.
@@
Figure 5 shows the time evolution of the total accretion stress $\alpha=(\langle{\rho u_{x}u_{y}}\rangle-\langle{B_{x}B_{y}}\rangle)/\langle{P_{\rm th}}\rangle$, the midplane $\beta$ including fluctuations ($\beta_{\rm mid}\equiv 8\pi c_{s}^{2}\langle{\rho}\rangle_{\rm th}/\langle{B^{2}}\rangle_{\rm th}$), and the midplane $\overline{B_{y}}$ for the simulations shown in Fig. 2 as well as several other illustrative examples. Most importantly, we see a large “gap” between the low-$\beta$ and high-$\beta$ states for each quantity. As well as clearly illustrating the bimodality in $\beta$ between the two states, these show more concretely that (i) the low-$\beta$ state supports much higher accretion rates and (ii) the low-$\beta$ state retains much more of the midplane magnetic flux from the initial conditions, whereas nearly all of the flux ($\gtrsim 95\%$) is lost in the evolution towards the high-$\beta$ state. It also highlights how the $\beta=0.1$, $H_{\beta}=\sqrt{2}$ initial conditions ($\beta$0.1, $\beta$R0.1, $\beta$0.1-mr, and $\beta$nR0.1-mr) are sensitive to slight changes in resolution and initial conditions: $\beta$0.1 transitions to the high-$\beta$ state rather slowly, while $\beta$R0.1 stays at low $\beta$ due to the extra contribution from $B_{x}$ that initially strengthens $B_{y}$ ($B_{x}B_{y}<0$); $\beta$0.1-mr (at coarser resolution) evolves back to lower $\beta$ and higher $\alpha$ after initially evolving similarly to $\beta$0.1, but adding a very weak radial field with $B_{x}B_{y}>0$ such that $B_{y}$ is initially made weaker ($\beta$nR0.1-mr) causes it to transition to the high-$\beta$ state (note that this $B_{r}$ with $\beta_{x}=1000$, is weaker than that which occurs self-consistently in the low-$\beta$ saturated state). We also show two cases that start instead from the quasi-Lorentzian initial conditions of Eq. 5. Because these profiles are already magnetically dominated at large $z$, they evolve less violently over the initial phases. Nonetheless, the bimodal behavior persists, with the lower-$\beta$ initial condition maintaining the strongly magnetized state and $\overline{B_{y}}$ increasing, while the higher-$\beta$ initial condition loses nearly all of its midplane flux. [^6] At $336^{3}$ resolution and $H_{\rm th}=0.43$, the high-$\beta$ state only sometimes produces well-formed cyclic behavior. Given the coarse effective resolution of the thermal midplane, and previous results documenting the resolution and box-size dependence of MRI turbulence (e.g., Davis et al., 2010; Ryan et al., 2017), this is not unexpected.
@@
Figure 6: Spatial profile of the various components of the accretion stress in the low-$\beta$ state ($\beta$R0.1; top) and the high-$\beta$ state ($\beta$1; bottom). We split $\alpha(z)$ into $\alpha_{K}\equiv\overline{u_{x}u_{y}}$, $\alpha_{\delta B}=-\overline{\delta B_{x}\delta B_{y}}$, and $\alpha_{\overline{B}}=-{\overline{B_{x}}\,\overline{B_{y}}}$, each normalized by $\langle{\rho}\rangle_{\rm th}c_{s}^{2}$. Magnetic contributions are dominant, with the total stress following a similar profile to $P_{B}$ except right in the midplane.
@@ 5.2 Turbulence and transport at low $\beta$
The turbulence profiles in the low-$\beta$ state are shown in Figs. 6 and 7. Figure 6 shows ${\alpha}(z)$ in $\beta$R0.1 and $\beta$1, splitting the total into its contributions from velocity fluctuations $\alpha_{K}=\overline{\rho u_{x}u_{y}}/\langle{P_{\rm th}}\rangle_{\rm th}$, magnetic fluctuations $\alpha_{\delta B}=-\overline{\delta B_{x}\delta B_{y}}/\langle{P_{\rm th}}\rangle_{\rm th}$, and the mean magnetic field $\alpha_{\overline{B}}=-\overline{B_{x}}\,\overline{B_{y}}/\langle{P_{\rm th}}\rangle_{\rm th}$ (note that the mean shear flow has been subtracted from $u_{y}$ before computing $\alpha_{K}$). The total ${\alpha}(z)$ profile is similar to that of the equilibrium-field $P_{B}$ (dotted line), aside from in the midplane, where there is a peak in the turbulent contributions and a drop off in the mean contribution. We interpret this change within $|z|\lesssim H_{\rm th}$ as a result of the MRI becoming dominant, whereas the Parker instability prevails at larger $z$. This interpretation is supported by the similar ratios between the $\alpha(z)$ components in the best-resolved high-$\beta$ simulation ($\beta$1), where the MRI remains active across a broad midplane region (cf. lower panel). We see that both MRI- and Parker-driven turbulence enable accretion ($\alpha>0$), but the MRI-driven midplane regions have a stronger relative contribution from magnetic, compared to velocity fluctuations.
@@
The thin black line in both panels illustrates a simple model of the turbulence, based on driving by a combination of the MRI and Parker instability (see also Johansen & Levin 2008). For both instabilities, we apply the standard reasoning that the turbulence saturates when the nonlinear turnover time $\tau_{\rm nl}\sim\ell/\delta u$ at the outer scale $\ell$ balances the instability’s growth rate. In regions far above the midplane, where the profiles of $B_{y}$, $\rho$, and $v_{\rm A}$ are all power laws, it is most natural to take $\ell\propto z$, such that the turbulence remains self-similar with $z$. Another possible choice, the lengthscale of the fastest growing mode, would scale in the same way for both the Parker and MRI: for Parker because it grows fastest at the largest available vertical scales (the global scales, $\propto z$), and for MRI because it grows fastest where $n\sim\Omega^{-1}v_{\rm A}/\ell\sim 1$ and $v_{\rm A}\propto z$ in such regions (see §3.1 and inset of Fig. 7). [^7] Similar arguments applied to the high-$\beta$ state would suggest that we should take $\ell\sim v_{\rm A}/\Omega$ of the fastest growing mode as the outer scale, where the $v_{\rm A}$ is that of the self-induced mean azimuthal field that can drive the MRI. This scale is very small, around $\ell\sim 4\times 10^{-2}H_{\rm th}$ for the $\beta$1 simulation (effectively the grid scale), in reasonable agreement with the morphology seen in Fig. 3. If $\ell$ is indeed grid limited, then since $\delta u\propto\ell$ (assuming nearly constant growth rate), we would expect the turbulence amplitude to decrease with finer resolution (complicated, perhaps, by the fact that $\overline{v_{\rm A}}$ may itself depend on $\delta u$). This is consistent with the results of Ryan et al. (2017) who found that the turbulence decreased its amplitude and scale with finer resolution in the high-$\beta$ state, even at extremely fine resolutions. In contrast, near the midplane, since the shape of $\rho$ is set by $H_{\rm th}$, we require $\ell\lesssim H_{\rm th}$. The Parker instability will not operate here because $B_{y}$ is nearly flat, while the scale at which the MRI growth rate peaks ($n\approx 1$ in Eq. 6 with $v_{\rm A}\sim c_{s}$) is $\ell\sim v_{\rm A}/\Omega\sim H_{\rm th}$, suggesting $\ell\sim H_{\rm th}$ is a reasonable choice (this is also plausibly consistent with the morphology seen in Fig. 3 for $\beta$R0.1).
@@
Figure 8: Testing different vertical boundary conditions, numerical options, and initial conditions (see table 1). As in Fig. 4, solid lines show $P_{B}$ and dashed lines show $P_{\rm th}$.The inset zooms into the midplane region.